OpenStudy (anonymous):

The area of triangle with base, b and height h, is given by A=1/2 BH If the area of a certain triangle is 65 square inches and the height is 26 inches, what is the base? Please show all of your work

OpenStudy (anonymous):

A=1/2bh A=65 h=26 b=? 65=1/2*b*26 65=13b b=65/13 b=5